Three charges q1=5, q2=4 and q3=3 are located at 0.6i, 0.3i, +0.8j respectively. Find the net force exerted on q due to the charges q2 and q3
1
Expert's answer
2021-10-21T21:41:01-0400

"F_{21}=\frac{kq_1q_2}{r^3}r"

"F_{21}=\frac{9\times10^9\times5\times4}{0.3^3}(-0.3\hat{i})"

"F_{21}=-2\times10^{12}\hat{i}N"

"F_{31}=\frac{kq_1q_3}{r^3}r"

"F_{21}=\frac{9\times10^9\times5\times3}{1^3}(-0.6\hat{i}+0.8\hat{j})"

"F_{21}=1.8\times10^{11}(-0.6\hat{i}+0.8\hat{j})N"